Can you please tell me if i have done this correctly: 86% of all U.S. families had incomes below the poverty level. Out of 400 randomly selected families whose householder had at least a Bachelores, 9 had incomes below the poverty level. Do the data provide sufficient evidence to conclude that the percentage of families that earned incomes below the poverty level was lower among those whoe householders had at least a Bachelors than among all U.S. Families? Usa a = 0.01
I did:
9/400=0.0225
z=0.0225-0.01/sqrt(0.01)(1-0.01)/400=0.022253, p = -0.4911
Can you please tell me if I am correct? Thank you for your help.

Answer by stanbon(75887)

86% of all U.S. families had incomes below the poverty level.
-----
Comment: That seems like an outlandish statement.
--------



Out of 400 randomly selected families whose householder had at least a Bachelores, 9 had incomes below the poverty level. Do the data provide sufficient evidence to conclude that the percentage of families that earned incomes below the poverty level was lower among those whoe householders had at least a Bachelors than among all U.S. Families? Usa a = 0.01
I did:
9/400=0.0225
z=0.0225-0.01/sqrt(0.01)(1-0.01)/400=0.022253, p = -0.4911
----
Comment: You do not use 0.01 in that way.
-------------------------------
Can you please tell me if I am correct?
---------------
Comment: If you can straighten out the percent of the general
population below the poverty line, you may be able to get the
correct answer.
As it is you are comparing 86% to 2.25%.
Hardly a close difference.
Cheers,
Stan H.
--------------
Ho: p = 0.86
Ha: p < 0.86
---
p-hat = 9/400 =